current through a conductor is I = (2t + 5)A. find the average current through a cross-section in the interval 0 to 4 s.
1
Expert's answer
2012-03-01T10:51:58-0500
I(t0) = I(0) = 2*0 + 5 = 5A;
I(t1) = I(4) = 2*4 + 5 = 13A;
I_average = (I(t0) + I(t1))/2 = (5+13)/2 = 18/2 = 9A.